OpenStudy (anonymous):

find the indicated outputs f(x=5xsquare 2- 2x f(2)=

OpenStudy (anonymous):

you should write f(x) = 5x^2 - 2x because it looks like you have a squared and a 2 instead

OpenStudy (anonymous):

If that's correct, then f(2) means everywhere you see x, replace it with 2: f(2) = 5(2)^2 - 2(2) = 5(4) - 4 = 20 - 4 = 16

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

I wasn't clear how to write it. thank u for helping me I really apperciate it